The 1.2 million preferred shares of Mighty Machines Ltd., which pay a dividend rate of 6.5 percent on a stated value of $40, are currently worth $42,000,000. What is the risk premium associated with these preferred shares if the risk-free rate is 4.25 percent?
1
Expert's answer
2019-11-13T09:02:23-0500

Rp = D/P + g - Rf = 40×0.065/(42,000,000/1,200,000) - 0.0425 = 0.0318 or 3.18%.